# The Hidden Power of Normalization: Exponential Capacity Control in Deep Neural Networks

### v1

@@ Abstract
Normalization methods are fundamental components of modern deep neural networks (DNNs). Empirically, they are known to stabilize optimization dynamics and improve generalization. However, the underlying theoretical mechanism by which normalization contributes to both optimization and generalization remains largely unexplained, especially when using many normalization layers in a DNN architecture.
In this work, we develop a theoretical framework that elucidates the role of normalization through the lens of capacity control. We prove that an unnormalized DNN can exhibit exponentially large Lipschitz constants with respect to either its parameters or inputs, implying excessive functional capacity and potential overfitting. Such bad DNNs are uncountably many. In contrast, the insertion of normalization layers provably can reduce the Lipschitz constant at an exponential rate in the number of normalization operations. This exponential reduction yields two fundamental consequences: (1) it smooths the loss landscape at an exponential rate, facilitating faster and more stable optimization; and (2) it constrains the effective capacity of the network, thereby enhancing generalization guarantees on unseen data. Our results thus offer a principled explanation for the empirical success of normalization methods in deep learning.
@@ 1 Introduction
Normalization techniques such as Batch Normalization ($\mathsf{BN}$}(Ioffe and Szegedy, 2015), Layer Normalization ($\mathsf{LN}$}(Ba et al., 2016), Group Normalization ($\mathsf{GN}$}(Wu and He, 2020), and Instance Normalization ($\mathsf{IN}$}(Ulyanov et al., 2016) have become indispensable components of modern deep neural networks (DNNs). In practice, they consistently enhance performance across a wide range of tasks. A classical motivation for normalization is to mitigate the internal covariate shift during training (Ioffe and Szegedy, 2015). Without normalization, the variance of inputs at intermediate layers may grow substantially (Ioffe and Szegedy, 2015; Bjorck et al., 2018)—even exponentially with depth (De and Smith, 2020; Zhang et al., 2019; Shao et al., 2020)—causing unstable gradients and slow convergence. Normalization counteracts this effect by rescaling inputs, thereby improving optimization stability.
Beyond optimization stability, normalization has long been observed to improve the generalization of DNNs, yet its underlying theoretical mechanism remains unclear. Several hypotheses have been proposed. Santurkar et al. (2018) showed that $\mathsf{BN}$} smooths the loss landscape by reducing the Lipschitz constant of the loss, which facilitates training. The ability to employ larger learning rates (Bjorck et al., 2018; Zhang et al., 2019) and the implicit regularization behavior observed in shallow architectures (Luo et al., 2019) further suggest a regularizing effect. However, a precise theoretical characterization of this regularization for deep networks is still missing (Wu et al., 2020). This leads to a long-standing open question:
Why do normalization methods help DNNs generalize well on unseen data?
In this work, we answer this question by showing that normalization methods provide explicit control over the functional capacity of a network, as quantified by its Lipschitz constant. Under the widely accepted assumption of large input variance, we demonstrate that normalization operations reduce the Lipschitz constant at an exponential rate with respect to normalization operations. This exponential reduction simultaneously smooths the optimization landscape and regularizes the network’s capacity, leading to improved generalization and faster convergence. Our analysis thus establishes a unified theoretical foundation for the empirical success of normalization methods.
@@
  • Lipschitz-based capacity analysis: We present a comprehensive theoretical analysis of normalization methods through the lens of Lipschitz-based capacity control.
  • Unnormalized networks exhibit extreme Lipschitz behavior: We show that a feedforward ReLU network (FFN) without normalization can have exponentially large or exponentially small Lipschitz constants (with respect to either weights or inputs), and that such configurations are uncountably many. Such constants exponentially depend on the depth. A huge Lipschitz constant w.r.t. the input implies that the network has an excessive funtional capacity and hence easily gets overfitting. In contrast, extremely small Lipschitz constant implies that the network is prone to underfiting. Therefore, within one architecture, there exist uncountably many models which potentially are either overfitting or underfitting.
@@
  • Normalization induces exponential capacity reduction: Under large-variance conditions, each normalization operation ($\mathsf{BN}$}, $\mathsf{LN}$}, $\mathsf{GN}$}, or $\mathsf{IN}$}) reduces the Lipschitz constant by a factor inversely proportional to the input variance. Multiple normalization layers thus lead to an exponential reduction, providing an explicit theoretical account of their regularization role.
@@
Overall, this work goes beyond simplified or shallow settings in prior studies to provide a rigorous theoretical understanding of normalization in modern deep architectures. While our analyses focus on feedforward networks, the revealed principles are general and can naturally extend to other architectures such as convolutional and transformer-based networks.
@@ 2.2 Notations
A bold character (e.g., $\bm{x}$}) often denotes a vector, while a bold big symbol (e.g., $\bm{D}$}) often denotes a matrix or set. Denote $\|\cdot\|$ as the max norm. Given ${\bm{x}}=(x_{1},...,x_{n})\in\mathbb{R}^{n}$ and a constant $a$, we denote ${\bm{x}}-a=(x_{1}-a,...,x_{n}-a)$. Later we often write ${\bm{A}}{\bm{x}}$ which is the multiplication of a matrix $\bm{A}$} with vector $\bm{x}$} organized in a suitable column. We denote $\epsilon$ as the (small) smoothing constant for some normalizers. ${\bm{x}}_{S}$ denotes a subset of $\bm{x}$}, corresponding to an index set ${\bm{S}}\subseteq[n]$. $|{\bm{S}}|$ denotes the size/cardinality of $\bm{S}$}.
A function ${\bm{y}}:({\mathcal{X}},d_{x})\rightarrow({\mathcal{Y}},d_{y})$ is said to be $L$-Lipschitz continuous if $d_{y}({\bm{y}}({\bm{x}}),{\bm{y}}({\bm{x}}^{\prime}))\leq Ld_{x}({\bm{x}},{\bm{x}}^{\prime})$ for any ${\bm{x}},{\bm{x}}^{\prime}\in{\mathcal{X}}$, where $d_{x}$ is a metric on $\mathcal{X}$}, $d_{y}$ is a metric on $\mathcal{Y}$}, and $L\geq 0$ is the Lipschitz constant. For simplicity we will consider both metrics to be max norm, i.e., $d_{x}({\bm{x}}_{1},{\bm{x}}_{2})=\|{\bm{x}}_{1}-{\bm{x}}_{2}\|$ and $d_{x}({\bm{y}}_{1},{\bm{y}}_{2})=\|{\bm{y}}_{1}-{\bm{y}}_{2}\|$. Denote $\|{\bm{y}}\|_{Lip}$ as the Lipschitz constant with respect to its input. For a function ${\bm{h}}({\bm{W}},{\bm{x}})$ depending on two variables, sometimes we write $\|{\bm{h}},{\bm{W}}\|_{Lip}$ to denote the Lipschitz constant of $\bm{h}$} w.r.t. $\bm{W}$} for a fixed $\bm{x}$}. Similarly, sometimes we can use $\|{\bm{h}},{\bm{x}}\|_{Lip}$, for a fixed $\bm{W}$}. For a linear operator ${\bm{A}}{\bm{x}}$, its Lipschitz constant is $\|{\bm{A}}{\bm{x}}\|_{Lip}=\|{\bm{A}}\|=\max_{1\leq k\leq n}\sum_{j\in[n]}|A_{kj}|$ which is the induced infinity norm of matrix $\bm{A}$}.
@@ 3.1 Batch normalization
Given $\epsilon>0$, let ${\bm{x}}=(x_{1},...,x_{n})$ be an input and ${\mathsf{BN}}({\bm{x}},\epsilon)$ be the normalization of $\bm{x}$}, where each input $x_{k}$ with population variance $\sigma_{k}^{2}$ is normalized as (1). Then $\|{\mathsf{BN}}\|_{Lip}=\|{1}/{\bm{\sigma}}\|$, where ${1}/{\bm{\sigma}}=\left({1}/{\sqrt{\sigma_{1}^{2}+\epsilon}},...,{1}/{\sqrt{\sigma_{n}^{2}+\epsilon}}\right)$.
@@ 3.2 Layer normalization
Consider ${\bm{x}}\in\mathbb{R}^{n}$ and $\sigma_{n}({\bm{x}})$ defined as (3). Given $\epsilon>0$, let $\sigma=\min_{{\bm{x}}\in\mathbb{R}^{n}}\sqrt{\sigma_{n}^{2}({\bm{x}})+\epsilon}$. We have $\|{\mathsf{LN}}\|_{Lip}\leq\left(1-\frac{1}{n}\right)/\sigma$.
Note the appearance of $1-\frac{1}{n}$ in the upper bound for $\|{\mathsf{LN}}\|_{Lip}$. This factor reveals an implicit penalty from $\mathsf{LN}$}. A larger $n$ will result in smaller penalty. In other words, for DNNs with $\mathsf{LN}$}, a smaller layer may be penalized stronger than a larger layer. Clearly such a penalty is imposed in an uncontrolled manner. On the other hand, this factor does not appear in Lemma 1 for $\mathsf{BN}$}, suggesting that $\mathsf{LN}$} may pose a stronger penalty than $\mathsf{BN}$}.
@@ 3.3 Group normalization
Let $\bigcup_{k=1}^{C}{\bm{S}}_{k}$ be a decomposition of $\{1,...,n\}$ into $C$ disjoint subsets, each of size at least 2. Assume that those subsets are used by $\mathsf{GN}$} to normalize an input ${\bm{x}}\in\mathbb{R}^{n}$. Denote $\sigma_{k}^{2}({\bm{x}})$ as the variance to compute (4) for group ${\bm{S}}_{k}$, and $\sigma_{k}^{2}=\min_{{\bm{x}}\in\mathbb{R}^{n}}\sigma_{k}^{2}({\bm{x}})$.
1. Denoting ${\bm{v}}=(v_{1},...,v_{C})$ where $v_{k}=\frac{1}{\sqrt{\sigma_{k}^{2}+\epsilon}}(1-\frac{1}{|{\bm{S}}_{k}|})$, we have $\|{\mathsf{GN}}\|_{Lip}\leq\|{\bm{v}}\|$.
2. Denoting $n_{g}=\max_{k}|{\bm{S}}_{k}|$ and $\sigma=\min_{k}\sqrt{\sigma^{2}_{k}+\epsilon}$, we have $\|{\mathsf{GN}}\|_{Lip}\leq\left(1-{1}/{n_{g}}\right)/\sigma$.
Comparing with Lemma 2, the factor $1-{1}/{n_{g}}$ in Lemma 3 suggests a stronger penalty due to $n_{g}\leq n$. When the groups in $\mathsf{GN}$} get smaller, a stronger penalty could be implicitly imposed. Note that such a penalty comes from the group size, which is entirely different from the motivation of large variances.
@@ 4.1 The Lipshitz constant of a DNN
These simple results lead to several important implications:
@@ 4.2 Smoothening DNNs by normalization
Let fixed activation functions $(g_{1},\dots,g_{K})$, whose Lipschitz constants are at most 1, and the bounds $(s_{1},\dots,s_{K})$. Let ${\bm{u}}_{i}=g_{i}({\mathsf{NO}}_{i}({\bm{W}}_{i}{\bm{u}}_{i-1}))$, with ${\bm{u}}_{0}={\bm{x}}$, be the neural network associated with weight matrices $({\bm{W}}_{1},\dots,{\bm{W}}_{i})$, where ${\mathsf{NO}}_{i}$ denotes a normalization operator at layer $i$. Let ${\mathcal{H}}_{no}=\{{\bm{u}}_{K}({\bm{x}}):{\bm{x}}\in\mathbb{R}^{n},\|{\bm{W}}_{i}\|\leq s_{i},\forall i\leq K\}$ be the family of normalized DNNs with $K$ layers.
@@
2. $\left\|{\bm{h}}_{no},{\bm{x}}\right\|_{Lip}\leq\sigma^{-K}P_{w}({\bm{h}}_{no})$, where $\sigma=-\max_{k}\|-\bm{\sigma}_{k}\|$.
@@
Consider any ${\bm{h}}_{no}\in{\mathcal{H}}_{no}$ in Definition 8 with operator ${\mathsf{NO}}_{k}(\cdot)\equiv{\mathsf{LN}}(\cdot,\epsilon)$ at any layer $k\leq K$. Denote $n_{k}$ as the size of the input and $\sigma_{k}$ as input variance before ${\mathsf{NO}}_{k}$, defined in Lemma 2. Then $\left\|{\bm{h}}_{no},{\bm{x}}\right\|_{Lip}\leq P_{w}({\bm{h}}_{no})\prod_{k=1}^{K}(1-\frac{1}{n_{k}})/{\sigma_{k}}$.
@@
Potentially exponential reduction due to normalization: When such large input variances occur, normalization layers can dramatically reduce the network’s effective Lipschitz constant. According to Corollary 10, $\mathsf{BN}$} yields a reduction factor of order $O\!\left(\prod_{k=1}^{K}\|1/{\bm{\sigma}_{k}}\|\right)$, while Corollary 11 indicates that $\mathsf{LN}$} provides a reduction of order $O\!\left(\prod_{k=1}^{K}(1-\frac{1}{n_{k}})/{\sigma_{k}}\right)$. As shown in Figure 2, the values of $\sigma_{k}$ are typically much greater than 1 after training. Consequently, the cumulative reduction factors can be extremely large in magnitude—potentially exponential in the number of normalization layers. For instance, in ResNet18 with 20 $\mathsf{BN}$} layers, we empirically observe $\prod_{k=1}^{K}1/{\sigma_{k}}\approx 10^{-25}$. Such a reduction is remarkably substantial and highlights the profound stabilizing effect of normalization on deep networks.

### v2

# The Hidden Power of Normalization Layers in Neural Networks: Exponential Capacity Control

@@ Abstract
Normalization layers are critical components of modern AI systems, such as ChatGPT, Gemini, DeepSeek, etc. Empirically, they are known to stabilize training dynamics and improve generalization ability. However, the underlying theoretical mechanism by which normalization layers contribute to both optimization and generalization remains largely unexplained, especially when using many normalization layers in a deep neural network (DNN).
In this work, we develop a theoretical framework that elucidates the role of normalization through the lens of capacity control. We prove that an unnormalized DNN can exhibit exponentially large Lipschitz constants with respect to either its parameters or inputs, implying excessive functional capacity and potential overfitting. Such bad DNNs are uncountably many. In contrast, the insertion of normalization layers provably can reduce the Lipschitz constant at an exponential rate in the number of normalization layers. This exponential reduction yields two fundamental consequences: (1) it smooths the loss landscape at an exponential rate, facilitating faster and more stable optimization; and (2) it constrains the effective capacity of the network, thereby enhancing generalization guarantees on unseen data. Our results thus offer a principled explanation for the empirical success of normalization methods in deep learning.
@@ 1 Introduction
Normalization methods such as Batch Normalization ($\mathsf{BN}$}(Ioffe and Szegedy, 2015), Layer Normalization ($\mathsf{LN}$}(Ba et al., 2016), Group Normalization ($\mathsf{GN}$}(Wu and He, 2020), Instance Normalization ($\mathsf{IN}$}(Ulyanov et al., 2016), and $\mathsf{RMSnorm}$} (Zhang and Sennrich, 2019) have become critical components of modern AI systems, such as ChatGPT, Gemini, DeepSeek, etc. Among them, $\mathsf{LN}$} or $\mathsf{RMSnorm}$} are frequently used in large language models (OpenAI et al., 2024; Abdin et al., 2024; Gu and Dao, 2024; Behrouz et al., 2025) and vision models (Peebles and Xie, 2023; Ma et al., 2024; Frans et al., 2025).
In practice, normalization layers consistently enhance performance of deep neural networks (DNNs) across a wide range of tasks. A classical motivation for normalization is to mitigate the internal covariate shift during training (Ioffe and Szegedy, 2015). Without normalization, the variance of inputs at intermediate layers may grow substantially (Ioffe and Szegedy, 2015; Bjorck et al., 2018)—even exponentially with depth (De and Smith, 2020; Zhang et al., 2019; Shao et al., 2020)—causing unstable gradients and slow convergence. Normalization counteracts this effect by rescaling inputs to improve optimization stability.
Beyond optimization stability, normalization layers has long been observed to improve the generalization of DNNs, yet its underlying theoretical mechanism remains unclear. Several hypotheses have been proposed. Santurkar et al. (2018) showed that $\mathsf{BN}$} smooths the loss landscape by reducing the Lipschitz constant of the loss, which facilitates training. The ability to employ larger learning rates (Bjorck et al., 2018; Zhang et al., 2019) and the implicit regularization behavior observed in shallow architectures (Luo et al., 2019) further suggest a regularizing effect. However, a precise theoretical characterization of this regularization for deep networks is still missing (Wu et al., 2020). This leads to a long-standing open question:
Why do normalization layers help DNNs generalize well on unseen data?
In this work, we answer this question by showing that normalization layers provide explicit control over the functional capacity of a network, as quantified by its Lipschitz constant. Under the widely accepted assumption of large input variance, we demonstrate that normalization operations reduce the Lipschitz constant at an exponential rate with respect to normalization operations. This exponential reduction simultaneously smooths the optimization landscape and regularizes the network’s capacity, leading to improved generalization and faster convergence. Our analysis thus establishes a unified theoretical foundation for the empirical success of normalization methods.
@@
  • Lipschitz-based capacity analysis: We present a comprehensive analysis of normalization methods through the lens of Lipschitz-based capacity control.
  • Unnormalized networks exhibit extreme Lipschitz behavior: We show that a feedforward ReLU network (FFN) without normalization can have exponentially large or exponentially small Lipschitz constants (with respect to either weights or inputs), and that such configurations are uncountably many. Such constants exponentially depend on the depth. A huge Lipschitz constant w.r.t. the input implies that the network has an excessive functional capacity and hence easily gets overfitting. In contrast, extremely small Lipschitz constant implies that the network is prone to underfitting. Therefore, within one architecture, there exist uncountably many models which potentially are either overfitting or underfitting.
@@
  • Normalization induces exponential capacity reduction: Under large-variance conditions, each normalization layer ($\mathsf{BN}$}, $\mathsf{LN}$}, $\mathsf{GN}$}, or $\mathsf{IN}$}) reduces the Lipschitz constant by a factor inversely proportional to the input variance. Multiple normalization layers thus lead to an exponential reduction, providing an explicit theoretical account of their regularization role.
@@
Overall, this work goes beyond simplified or shallow settings in prior studies to provide a rigorous theoretical understanding of normalization layers in modern deep architectures. While our analyses focus on feedforward networks, the revealed principles are general and can naturally extend to other architectures such as convolutional and transformer-based networks.
@@ 2.2 Notations
A bold character (e.g., $\bm{x}$}) often denotes a vector, while a bold big symbol (e.g., $\bm{D}$}) often denotes a matrix or set. Denote $\|\cdot\|$ as the $\ell_{2}$ norm. Given ${\bm{x}}=(x_{1},...,x_{n})\in\mathbb{R}^{n}$ and a constant $a$, we denote ${\bm{x}}-a=(x_{1}-a,...,x_{n}-a)$. Later we often write ${\bm{A}}{\bm{x}}$ which is the multiplication of a matrix $\bm{A}$} with vector $\bm{x}$} organized in a suitable column. We denote $\epsilon$ as the (small) smoothing constant for some normalizers. ${\bm{x}}_{S}$ denotes a subset of $\bm{x}$}, corresponding to an index set ${\bm{S}}\subseteq[n]$. $|{\bm{S}}|$ denotes the size/cardinality of $\bm{S}$}.
A function ${\bm{y}}:({\mathcal{X}},d_{x})\rightarrow({\mathcal{Y}},d_{y})$ is said to be $L$-Lipschitz continuous if $d_{y}({\bm{y}}({\bm{x}}),{\bm{y}}({\bm{x}}^{\prime}))\leq Ld_{x}({\bm{x}},{\bm{x}}^{\prime})$ for any ${\bm{x}},{\bm{x}}^{\prime}\in{\mathcal{X}}$, where $d_{x}$ is a metric on $\mathcal{X}$}, $d_{y}$ is a metric on $\mathcal{Y}$}, and $L\geq 0$ is the Lipschitz constant. For simplicity we will consider both metrics to be $\ell_{2}$ distance, i.e., $d_{x}({\bm{x}}_{1},{\bm{x}}_{2})=\|{\bm{x}}_{1}-{\bm{x}}_{2}\|$ and $d_{x}({\bm{y}}_{1},{\bm{y}}_{2})=\|{\bm{y}}_{1}-{\bm{y}}_{2}\|$. Denote $\|{\bm{y}}\|_{Lip}$ as the Lipschitz constant with respect to its input. For a function ${\bm{h}}({\bm{W}},{\bm{x}})$ depending on two variables, sometimes we write $\|{\bm{h}},{\bm{W}}\|_{Lip}$ to denote the Lipschitz constant of $\bm{h}$} w.r.t. $\bm{W}$} for a fixed $\bm{x}$}. Similarly, sometimes we can use $\|{\bm{h}},{\bm{x}}\|_{Lip}$, for a fixed $\bm{W}$}. For a linear operator ${\bm{A}}{\bm{x}}$, its Lipschitz constant is $\|{\bm{A}}{\bm{x}}\|_{Lip}=\|{\bm{A}}\|$ which is the spectral norm of matrix $\bm{A}$}.
@@ 3.1 Batch normalization
Given $\epsilon>0$, let ${\bm{x}}=(x_{1},...,x_{n})$ be an input and ${\mathsf{BN}}({\bm{x}},\epsilon)$ be the normalization of $\bm{x}$}, where each input $x_{k}$ with population variance $\sigma_{k}^{2}$ is normalized as (1). Then $\|{\mathsf{BN}}\|_{Lip}=\|{1}/{\bm{\sigma}}\|$, where ${1}/{\bm{\sigma}}=\left((\sigma_{1}^{2}+\epsilon)^{-0.5},...,(\sigma_{n}^{2}+\epsilon)^{-0.5}\right)$.
@@ 3.2 Layer normalization
Consider any ${\bm{x}}\in\mathbb{R}^{n}$ and $\sigma_{n}$ defined as (3). For any $\epsilon>0$, we have $\left\|\frac{\partial{\mathsf{LN}}({\bm{x}},\epsilon)}{\partial{\bm{x}}}\right\|\leq\frac{1}{\sqrt{\sigma^{2}_{n}+\epsilon}}$. Denoting $\sigma=\min_{{\bm{x}}\in\mathbb{R}^{n}}\sqrt{\sigma_{n}^{2}+\epsilon}$, we further have $\|{\mathsf{LN}}\|_{Lip}\leq 1/\sigma$.
Lemma 2 shows that the sensitivity of $\mathsf{LN}$} is entirely governed by its variance normalization term. Geometrically, $\mathsf{LN}$} first projects the input onto the mean-zero subspace and then rescales it by the inverse empirical standard deviation. The projection step is non-expansive (spectral norm equal to 1), so any potential amplification arises solely from the scaling factor $(\sigma_{n}^{2}+\epsilon)^{-0.5}$. The lemma therefore formalizes a clean operator-theoretic statement: $\mathsf{LN}$} cannot expand perturbations by more than $(\sigma_{n}^{2}+\epsilon)^{-0.5}$. In particular, the stabilizer $\epsilon$ ensures global Lipschitz continuity, preventing gradient blow-up when the input variance approaches zero.
Compared with $\mathsf{BN}$}, this behavior is structurally simpler and more localized. $\mathsf{LN}$} operates independently on each sample, so its Jacobian and Lipschitz constant depend only on that sample’s feature variance. In contrast, $\mathsf{BN}$} empirically estimates the mean and variance of each input from a batch and hence $\mathsf{BN}$} normalizes across the batch, making its Jacobian couple different examples and rendering its sensitivity dependent on batch statistics and inter-sample correlations. Consequently, $\mathsf{LN}$} admits a deterministic per-sample Lipschitz bound, while $\mathsf{BN}$}’s effective smoothness is inherently batch-dependent.
@@ 3.3 Group normalization
Let $\bigcup_{k=1}^{C}{\bm{S}}_{k}$ be a decomposition of $\{1,...,n\}$ into $C$ disjoint subsets, each of size at least 2. Assume that those subsets are used by $\mathsf{GN}$} to normalize an input ${\bm{x}}\in\mathbb{R}^{n}$. Denote $\sigma_{k}^{2}({\bm{x}})$ as the variance to compute (4) for group ${\bm{S}}_{k}$, and $\sigma_{k}^{2}=\min_{{\bm{x}}\in\mathbb{R}^{n}}\sigma_{k}^{2}({\bm{x}})$. We have $\|{\mathsf{GN}}\|_{Lip}\leq\sum_{k=1}^{C}(\sigma_{k}^{2}+\epsilon)^{-0.5}$.
@@ 4.1 The Lipschitz constant of a DNN
These results lead to several implications:
@@ 4.2 Smoothening DNNs by normalization
Let fixed activation functions $(g_{1},\dots,g_{K})$, whose Lipschitz constants are at most 1, and the bounds $(s_{1},\dots,s_{K})$. Let ${\bm{h}}_{i}=g_{i}({\mathsf{NO}}_{i}({\bm{W}}_{i}{\bm{h}}_{i-1}))$, with ${\bm{h}}_{0}={\bm{x}}$, be the neural network associated with weight matrices $({\bm{W}}_{1},\dots,{\bm{W}}_{i})$, where ${\mathsf{NO}}_{i}$ denotes a normalization operator at layer $i$. Let ${\mathcal{H}}_{no}=\{{\bm{h}}_{K}({\bm{x}}):{\bm{x}}\in\mathbb{R}^{n},\|{\bm{W}}_{i}\|\leq s_{i},\forall i\leq K\}$ be the family of normalized DNNs with $K$ layers.
@@
2. $\left\|{\bm{h}}_{no},{\bm{x}}\right\|_{Lip}\leq\sigma^{-K}P_{w}({\bm{h}}_{no})$, where $\sigma=\min_{k}\|1/\bm{\sigma}_{k}\|^{-1}$.
@@
Consider any ${\bm{h}}_{no}\in{\mathcal{H}}_{no}$ in Definition 8 with operator ${\mathsf{NO}}_{k}(\cdot)\equiv{\mathsf{LN}}(\cdot,\epsilon)$ at any layer $k\leq K$. Denote $n_{k}$ as the size of the input and $\sigma_{k}$ as input variance before ${\mathsf{NO}}_{k}$, defined in Lemma 2. Then $\left\|{\bm{h}}_{no},{\bm{x}}\right\|_{Lip}\leq P_{w}({\bm{h}}_{no})\prod_{k=1}^{K}\sigma_{k}^{-1}$.
@@
Potentially exponential reduction due to normalization: When such large input variances occur, normalization layers can dramatically reduce the network’s effective Lipschitz constant. According to Corollary 10, $\mathsf{BN}$} yields a reduction factor of order $O\!\left(\prod_{k=1}^{K}\|1/{\bm{\sigma}_{k}}\|\right)$, while Corollary 11 indicates that $\mathsf{LN}$} provides a reduction of order $O\!\left(\prod_{k=1}^{K}1/\sigma_{k}\right)$. As shown in Figure 2, the values of $\sigma_{k}$ are typically much greater than 1 after training. Consequently, the cumulative reduction factors can be extremely large in magnitude—potentially exponential in the number of normalization layers. For instance, in ResNet18 with 20 $\mathsf{BN}$} layers, we empirically observe $\prod_{k=1}^{K}1/{\sigma_{k}}\approx 10^{-25}$. Such a reduction is remarkably substantial and highlights the profound stabilizing effect of normalization on deep networks.